### v1

# Maximal information at the edge of stability in excitatory-inhibitory neural populations

Abstract
Understanding how the complex connectivity structure of the brain shapes its information-processing capabilities is a long-standing question. Here, by focusing on a paradigmatic architecture, we study how the neural activity of excitatory and inhibitory populations encodes information on external signals. We show that information is maximized at the edge of stability, where excitation is balanced by inhibition. When the input switches among different stimuli, this maximum corresponds to the entropy of the external switching dynamics. By analyzing the case of a prolonged stimulus, we find that stronger inhibition is needed to maximize the instantaneous sensitivity, revealing an intrinsic trade-off between short-time responses and long-time accuracy. In agreement with recent experimental findings, our results open the avenue for a complete information-theoretic understanding of how and why inhibition strength should be tuned to optimize information-processing capabilities.
preprint: APS/123-QED
From sensory perception to task-driven behaviors and decision-making processes, the brain constantly receives and integrates large amounts of environmental information. Both the encoding and processing of this information in the cortex involve a complex interplay among different neuronal populations koch1994large ; sakurai1996population ; quiroga2007decoding ; vyas2020computation . Understanding the basic mechanisms behind such information processing is thus a central topic in systems neuroscience. Several studies have investigated neural encoding at the level of individual neurons, showing that certain neurons selectively respond to specific features of incoming stimuli, such as spatial or temporal frequency, orientation, position, or depth kriegeskorte2021neural ; barlow1967neural ; campbell1968angular ; henry1974orientation , leading to a sparse encoding olshausen1996emergence . However, due to advancements in the ability to simultaneously record the activity of large numbers of neurons across various brain areas, recent decades have witnessed a shift of research focus towards the investigation of collective dynamics of neural populations gao2015simplicity ; buzsaki2004large ; brown2004multiple . Remarkably, the trajectories of such populations are typically constrained in low-dimensional manifolds in the high-dimensional space of neural activity kriegeskorte2021neural ; buonomano2009state ; pandarinath2018inferring , suggesting that the entire population dynamically encodes stimulus variables in this reduced cunningham2014dimensionality ; mante2013context ; gallego2017neural ; remington2018flexible or coarse-grained meshulam2019coarse ; nicoletti2020scaling neural state space. Tools from information theory have been used to measure the amount of information that the response of a neural system conveys on a stimulus brunel1998mutual ; quian2009extracting ; bernardi2015frequency ; panzeri2022structures , for instance through mutual or Fisher information shannon1948mathematical ; blahut1987principles . Yet, understanding how the emergent information properties depend on the underlying dynamics of the neural populations remains an open question.
Recurrent neural networks trained on cognitive tasks have shown that low-dimensional subspaces naturally emerge to support flexible computations at the population level dubreuil2022role ; beiran2023parametric . However, the internal connectivities of these artificial networks do not adhere to biological constraints such as Dale’s principle, which states that neurons must be either excitatory or inhibitory. This neuro-anatomical feature has profound implications for functionality, as it underpins amplification mechanisms murphy2009balanced ; hennequin2014optimal ; apicella2021emergence ; schaub2015emergence and efficient transmission baggio2020efficient .
It has been recently shown that strong recurrent coupling and inhibition stabilization are common features of the cortex sanzeni2020inhibition . Crucially, such a finely balanced state not only prevents instability but may enhance the system’s computational properties as well. Networks operating in this near-critical state often perform better in information processing tasks and complex computations in general langton1990computation ; bertschinger2004real ; beggs2008criticality , while exhibiting optimal sensitivity to sensory stimuli kinouchi2006optimal ; shew2009neuronal . Such an interplay between strong excitatory coupling and compensatory inhibition is shaped by the connectivity structure between neural populations, which makes theoretical studies particularly challenging.
In this work, we explicitly tackle the problem of quantifying the information encoded by neuronal subpopulations on an external stochastic stimulus. By leveraging neural dynamical models and focusing on the connectivity between excitatory and inhibitory populations, we are able to compute the information between the neuronal system and the stimulus in suitable regimes and provide analytical bounds on the mutual information. Specifically, we demonstrate that an excitatory-inhibitory balanced state is necessary not only to ensure stability but also for maximizing information encoding at the steady state. Further, by studying the response of the system to a single stochastic perturbation of varying intensity, we reveal an intrinsic trade-off between the optimal response at short and long times. In particular, the time-resolved form of the mutual information shows that global inhibition acts to regulate the total information encoded and the sensitivity of the system’s response.
To retain physical interpretability, we focus on a coarse-grained model of two neuronal subpopulations, one excitatory and one inhibitory, described by their respective firing rates $\rho_{E}$ and $\rho_{I}$. The excitatory population receives a time-varying external input $h(t)$, representing the stimuli the neurons seek to encode in their firing rate dynamics. We are interested in the fluctuations around the steady states of the two populations, $\rho_{\mu}^{*}$, which we define as $x_{\mu}(t)=\rho_{\mu}(t)-\rho_{\mu}^{*}$ with $\mu\in\{E,I\}$. This model has been widely used in the literature as it captures essential properties of neuronal connectivity murphy2009balanced ; schaub2015emergence ; christodoulou2022regimes ; wu2023inhibition . Following these previous works, we obtain the Langevin equations:
$\tau\frac{dx_{\mu}}{dt}=-r_{\mu}x_{\mu}+\sum_{\nu\in{E,I}}A_{\mu\nu}x_{\nu}+\sqrt{2D_{\mu}}\xi_{\mu}+h(t)\delta_{\mu,E}$ (1)
where $\tau$ is the characteristic timescale of neural dynamics, $r_{\mu}$ is the decay term of the activity, $\xi_{\mu}$ are independent white noises, and $A_{\mu\nu}$ is an element of the synaptic connectivity matrix, that is parametrized as
$\hat{A}=\begin{pmatrix}w&-kw\\ w&-kw\end{pmatrix}\,,$ (2)
with $w\geq 0$ and $k\geq 0$. Thus, $w$ measures the overall excitation strength, while $k$ quantifies the relative intensity of the inhibition. One can immediately show that, if $r_{\mu}=r$, the system is stable when $k>k_{c}=1-r/w$, while for $k<k_{c}$ the excitation is too strong for the inhibition to stabilize it.
To model changes in the external environment, we posit the input as a jump process between a ground state, $h_{0}$, and a set of $M$ discrete environmental states, $h_{i}$. These states might represent sensory stimuli of different intensities, behavioral states, or motor commands. For simplicity, we take these states to be equispaced, such that $h_{i}=h_{0}+i\Delta h$, with $h_{0}=0$ representing the absence of external signals. The environment can transition from this ground state to any other state with uniform transition rates, which we denote by $w_{0\to i}=w_{\uparrow}$. Likewise, the transition rate to the ground state from any other state is $w_{i\to 0}=w_{\downarrow}$. All other transition rates are set to zero, i.e., there is no direct switch from one stimulus to another. In this setting, all environmental states are equally likely and the neural populations must respond to the stochastic jumps among them. The characteristic timescale of this environmental process is $\tau_{\mathrm{input}}=(w_{\uparrow}+w_{\downarrow})^{-1}$. We sketch the model in Fig. 1a.
Figure 1: (a) Sketch of the model, describing a population of excitatory ($E$, green) and inhibitory neurons ($I$, blue) in a linearized regime evolving on a timescale $\tau$. An input $h$ stimulates activity in the excitatory population with a timescale $\tau_{\mathrm{input}}$. (b-c) If $\tau_{\mathrm{input}}\ll\tau$, the neural populations are not able to resolve the different inputs. In the opposite limit, the joint probability $p(E,I)$ displays instead peaks around the different input strengths. (d) The mutual information $I_{\mathbf{x},h}$ is zero in the fast-inputs regime, but sharply increases when $\tau_{\mathrm{input}}\gg\tau$, signaling that the neural populations are capturing information on the input in this regime. Simulations obtained with $M=2$, $\tilde{w}_{\uparrow}=1/3,\tilde{w}_{\downarrow}=2/3,D=1/2,r=1,\tau=1$.
We seek to understand how much information the neuronal network can capture on the external inputs at stationarity. To this end, we compute the mutual information shannon1948mathematical ; blahut1987principles
[…]
$=H_{\bm{x}}+H_{\mathrm{input}}-H_{\bm{x},\mathrm{input}}$ (3)
which is nothing but the Kullback-Leibler divergence $D_{\mathrm{KL}}$ between the joint steady-state probability of the inputs and the neural activity, $p^{\mathrm{st}}_{i,\bm{x}}$, with $\bm{x}=(x_{E},x_{I})$, and the corresponding marginal distributions, $p^{\mathrm{st}}_{\bm{x}}$ and $\pi_{i}^{\mathrm{st}}$ cover1999elements . As such, $I_{\bm{x},h}^{\mathrm{st}}$ quantifies how much $\bm{x}$ and $h$ are dependent on one another in terms of how much information they share in their joint probability distribution. As in Eq. (3), the mutual information can be rewritten in terms of the differential entropy of the excitatory and inhibitory populations, $H_{\bm{x}}$, the Shannon entropy of the external inputs, $H_{\mathrm{input}}$, and the joint entropy of the neural activity and environment, $H_{\bm{x},\mathrm{input}}$. Without loss of generality, we now assume that the populations have the same diffusion coefficient, i.e., $D_{\mu}=D$, and the same decay rate, i.e., $r_{\mu}=r$.
From Eq. (1), the joint probability $p_{i,\bm{x}}(t)$ is governed by the following Fokker-Planck equation:
$\partial_{t}p_{i,\bm{x}}(t)=$ $\dfrac{1}{\tau}\sum_{\mu=E,I}\biggl{[}\partial_{\mu}[(\tilde{F}_{i\mu}(\bm{x})p_{i,\bm{x}}(t)]+\partial_{\mu}^{2}p_{i,\bm{x}}(t)\biggl{]}+$
$+\dfrac{1}{\tau_{\mathrm{input}}}\sum_{j=0}^{M}\biggl{[}\tilde{w}_{j\to i}p_{j,\bm{x}}(t)-\tilde{w}_{i\to j}p_{i,\bm{x}}(t)\biggl{]}$ (4)
where $\tilde{w}_{j\to i}=\tau_{\mathrm{input}}w_{j\to i}$ are the rescaled transition rates, $\tilde{F}_{i\mu}(\bm{x})=-rx_{\mu}+\sum_{\nu}\mathcal{A}_{\mu\nu}x_{\nu}+h_{i(t)}\delta_{\mu,E}$ with $i(t)$ the environmental state at time $t$, and we used the shorthand notation $\partial_{x_{\mu}}:=\partial_{\mu}$. Finding an explicit solution of Eq. (4) is, in general, a formidably challenging task. However, exact solutions can be accomplished in a timescale separation limit nicoletti2021mutual ; nicoletti2022mutual ; nicoletti2022information .
We find that, in the limit of a fast-evolving input $\tau_{\mathrm{input}}\ll\tau$, the joint probability of the system factorizes as $p_{i,\bm{x}}^{\mathrm{st}}=p_{\bm{x}}^{\mathrm{st}}\pi_{i}^{\mathrm{st}}$ (see Supplemental Material suppinfo_ref ). In particular, $p_{\bm{x}}^{\mathrm{st}}$ corresponds to the solution of the Langevin equations in Eq. (1) with an effective input $\tilde{h}_{\mu}=\delta_{\mu,E}\sum_{i}h_{\mu,i}\pi_{i}^{\mathrm{st}}$ (Figure 1b). In this regime, we clearly have that the mutual information between the neural populations and the input vanishes, i.e., $I_{\bm{x},h}\to 0$ when $\tau_{\mathrm{input}}/\tau\to 0$. Indeed, the stationary solution of the system tells us that the neural activity is only influenced by the average input, as it cannot resolve its fast temporal evolution. On the other hand, in the limit of a slowly evolving external input $\tau_{\mathrm{input}}\gg\tau$, we find that the system is described by the stationary probability $p_{i,\bm{x}}(t)=p_{\bm{x}|i}^{\mathrm{st}}\pi_{i}(t)$, where $p^{\mathrm{st}}_{\bm{x}|i}$ is the probability of the excitatory and inhibitory populations conditioned on a given environmental state $h_{i}$. In particular, $p^{\mathrm{st}}_{\bm{x}|i}$ is the stationary solution of the system of Langevin equations in Eq. (1) with constant input $h_{i}$, which is a multivariate Gaussian distribution $\mathcal{N}(\bm{m}^{\mathrm{st}}_{i},\hat{\Sigma}^{\mathrm{st}})$. The mean $\bm{m}^{\mathrm{st}}_{i}$ is given by
$\begin{pmatrix}{m}^{\mathrm{st}}_{E,i}\\ {m}^{\mathrm{st}}_{I,i}\end{pmatrix}=\left(\hat{R}-\hat{A}\right)^{-1}\begin{pmatrix}h_{i}\\ 0\end{pmatrix}$
where $R_{\mu\nu}=r_{\mu}\delta_{\mu\nu}$, whereas the covariance $\hat{\Sigma}^{\mathrm{st}}$ satisfies the Lyapunov equation:
$\sum_{\nu\in\{E,I\}}\left[\left(A_{\alpha\nu}-R_{\alpha\nu}\right)\Sigma^{\mathrm{st}}_{\nu\beta}+\Sigma^{\mathrm{st}}_{\alpha\nu}\left(A_{\nu\beta}^{T}-R_{\nu\beta}\right)\right]=-\delta_{\alpha\beta}$
as we show in the Supplemental Material suppinfo_ref . It is worth noting here that, since the input acts as an additional drift, it only changes the average of the distribution with respect to the case of no input. Therefore, the stationary probability distribution of the neural populations is the Gaussian mixture $p_{\bm{x}}^{\mathrm{st}}=\sum_{i}\pi_{i}^{\mathrm{st}}p_{\bm{x}|i}^{\mathrm{st}}$. We show a typical trajectory of the system and the corresponding probability distribution of neural activity in Figure 1c.
Even though the entropy of a Gaussian mixture cannot be written in a closed form, by employing the bounds proposed in kolchinsky2017estimating , we obtain an upper and a lower bound on the mutual information starting from the Chernoff-$\alpha$ divergence and the Kullback-Leibler divergence between the mixture components (see Supplemental Material suppinfo_ref ). We have that $I^{(b)}(\eta/4)\leq I_{\bm{x},h}\leq I^{(b)}(\eta)$, where
$I^{(b)}(\eta)$ $=-\sum_{i=0}^{M}\pi_{i}\log[\sum_{j=0}^{M}\pi_{j}e^{-(j-i)^{2}\eta}\biggl{]}$ (5)
with
$\eta=\frac{\Delta h^{2}}{4Dr}\dfrac{(2r+(k-1)w)(2r^{2}+(3k-1)w+(k^{2}+1)w^{2})}{(r+(k-1)w)(2r^{2}+2(k-1)rw+(k^{2}+1)w^{2})}\;.$
Since $I^{(b)}(\eta/4)>0$, we have that $I_{\bm{x},h}$ is always non-zero. Eq. (5) shows that, in the limit of a slow input, the excitatory and inhibitory populations are able to capture information on the external stimulus. In the intermediate regime between the fast- and slow-input limits, we cannot solve the Fokker-Planck equation explicitly. However, a direct simulation of the system shows that the mutual information smoothly interpolates between the two regimes, as we see in Figure 1d. Taken together, our results underscore the significance of timescales for neuronal circuits and their capability of processing information on external time-varying stimuli butts2007temporal ; das2019critical ; mariani2021b ; nicoletti2024information .
Figure 2: (a) Mutual information between the neural activity $\bm{x}=(x_{E},x_{I})$ and the input $h$, as a function of the excitation strength $w$ and the relative intensity of inhibition $k$. If $k<k_{c}$ ($k_{c}$ is the black dotted line), the system is unstable. (b-c) Information between the neural activity and the inputs is maximized at the edge of stability when $k\to k_{c}$. In this limit, $I_{\bm{x},h}$ converges to the entropy of the input $H_{\mathrm{input}}$, which quantifies the information contained in the input.
Figure 3: (a) Dynamics of the mutual information in the presence of a prolonged stochastic input. (b-c) The stationary mutual information $I_{\bm{x},h}^{\mathrm{st}}$ diverges at the edge of stability, whereas the sensitivity $\chi_{\bm{x},h}=\partial_{t}^{2}I_{\bm{x},h}(t)|_{t=t_{\mathrm{stim}}}$ peaks at intermediate values of $k$ for large $w$. (d-e) The peak of the sensitivity occurs for $k_{\mathrm{max}}(w)>k_{c}$, while $I_{\bm{x},h}^{\mathrm{st}}\to+\infty$ for $k\to k_{c}$. (f) Thus, a larger inhibition strength $k$ benefits the response at short times in terms of how quickly information increases when the input signal arrives. On the contrary, at long times, information is maximized by reducing $k$ and approaching the edge of stability at $k=k_{c}$.
Crucially, the synaptic strengths of the excitatory and inhibitory populations drastically affect their mutual information with the input. Indeed, as we show in Figure 2a, $I_{\bm{x},h}$ strongly depends on the interplay between excitation and inhibition. Furthermore, the bounds in Eq. (5) tighten as $k$ approaches $k_{c}$, eventually collapsing to one single value in the limit $k\to k_{c}$, which corresponds to the edge of stability of the system (see Fig. 2b):
$I_{\bm{x},h}\underset{k\to k_{c}}{\longrightarrow}H_{\mathrm{input}}\;.$ (6)
Eq. (6) tells us that, at the edge of stability, the neural populations are able to fully capture the information contained in the external input, which is exactly its entropy $H_{\mathrm{input}}$. Intriguingly, we also find that this is the maximum value the mutual information can attain, as shown in Figure 2b-c. Thus, modulation of the inhibition strength plays a prime role in determining how efficiently the system can encode the external inputs, and the corresponding mutual information sharply increases as the edge of stability is approached.
So far, we have considered the steady-state response of the system to a time-varying input. However, understanding how neural populations dynamically acquire information when they receive a single stimulus, and how quickly they can do so, is crucial as well. To this end, we now introduce a scenario where the system in its stationary state is perturbed by an external stochastic input. In this case, the mutual information in time reads
[…]
where $p_{h}^{\mathrm{st}}$ is the distribution of the strength of the stimulus, and $H_{\bm{x}|h}$ is the conditional differential entropy of $\bm{x}$ at a given input value $h$. We assume that $p_{h}^{\mathrm{st}}$ is fully characterized by the mean input strength, $\mu_{h}$, and its variance, $\sigma_{h}$, so that $p_{h}=\mathcal{N}(\mu_{h},\sigma_{h})$. Then, we have that
$I_{\bm{x},h}(t)$ $=\dfrac{1}{2}\log\dfrac{\det\left[\Sigma^{\mathrm{st}}+K(t)\begin{pmatrix}\sigma_{h}^{2}&0\\ 0&0\end{pmatrix}K(t)^{T}\right]}{\det\left(\Sigma^{\mathrm{st}}\right)}$ (8)
where $\Sigma^{\mathrm{st}}$ is the covariance matrix of the unperturbed system at steady state, and $K(t)$ is the time-dependent gain matrix that we derive in the Supplemental Material suppinfo_ref . In Figure 3a, we plot the time evolution of the mutual information. At long times, we find once more that the mutual information is maximized at the edge of stability (Figure 3b), with $I_{\bm{x},h}$ diverging as $k\to k_{c}$. We note that, since the differential entropy for the continuous input distribution is not necessarily positive, the bounds in Eq. (5) cannot be straightforwardly applied. In particular, while the maximal information content of the input was associated with its switching dynamics in the previous setting, there is now no a priori limit to the information that the system can encode.
The scenario becomes more intricate at short times after the stimulus. In the inhibition-stabilized regime, where $w>1$, the response of the neural populations exhibits a faster increase for stronger excitatory couplings away from the edge of instability. To assess the system’s responsiveness, we introduce a metric of sensitivity defined as $\chi_{\bm{x},h}=\partial_{t}^{2}I_{\bm{x},h}(t)|_{t=t_{\mathrm{stim}}}$, which quantifies the instantaneous increase in mutual information when the stimulus arrives at time $t=t_{\mathrm{stim}}$ derivative_note . In Figure 3c, we show that $\chi_{\bm{x},h}$ peaks at an optimal inhibition strength $k_{\rm max}>1$, whose complete expression is given in suppinfo_ref . Crucially, this optimal value depends on the excitation strength as well, i.e., $k_{\rm max}\equiv k_{\rm max}(w)$ (see Fig. 3d). This highlights that the inhibition regime for the optimal response at short times is drastically different from that at long times (see Figs. 3d-e). In particular, since $k_{c}(w)<k_{\mathrm{max}}(w)$ for all $w$, our results unravel a fundamental trade-off between achieving maximum accuracy and the speed at which the neural populations encode information about the external stimulus, akin to a speed-accuracy trade-off lan2012energy . As we show in Figure 3f, long-time accuracy is generally achieved at lower inhibition, whereas sensitivity maximization requires a larger value of $k$.
Overall, our analysis marks a significant step towards the understanding of how the structure of connectivity shapes information encoding in neuronal population dynamics. We showed analytically in an exactly solvable regime that the mutual information between an input and the receiving neuronal populations is maximized at the edge of stability, highlighting the importance of the balance between excitation and inhibition. As non-normal synaptic interactions are crucial for achieving this optimal state, our findings confirm that the underlying structural connectivity is essential for the emergence of critical behaviors that support optimal information encoding and complex dynamical evolution both in whole-brain connectomes barzon2022criticality and in artificial recurrent neural networks kerg2019non . Notably, alterations in excitatory-inhibitory balance have been experimentally related to the loss of information-processing efficiency observed in pathological conditions dehghani2016dynamic ; sohal2019excitation . Further, our results point towards the hypothesis that the bran might be poised at criticality to optimize its information-processing capabilities beggs2008criticality ; o2022critical ; mariani2021b ; mariani2022disentangling , and that computation at edge-of-chaos may be optimal langton1990computation .
Our predictions are also consistent with recent experimental studies in which theoretical tools from response theory have been applied to extensive whole-brain neuronal recordings. The emergent dynamics of several brain regions has been shown to lie at the edge of stability dahmen2019second ; morales2023quasiuniversal , with a distance from instability that only slightly varied along the cortex. Such heterogeneity might be explained as an increase in the inhibition level wang2020macroscopic . Our findings suggest that this observed feature may be related to the tuning of sensitivity to different timescales murray2014hierarchy ; manea2022intrinsic . Importantly, the external stimulus considered here may be immediately generalized to a high-dimensional signal representing, for example, multiple stimuli with different characteristics (e.g., frequency, intensity) targeting spatially separated populations.
Although we focused on a paradigmatic - yet widely used - model, our approach applies in principle to any neuronal connectivities and can be used to investigate more detailed synaptic structures and their effects on information-processing performances. Although it will be crucial to understand how strongly non-linear regimes affect the behavior of the mutual information and the information-theoretic features of the system in general, our work paves the way to the unraveling of the fundamental mechanisms supporting information encoding and sensitivity in neuronal networks.
Acknowledgements.
G.N. acknowledges funding provided by the Swiss National Science Foundation through its Grant CRSII5_186422.

### v2

@@
-Maximal information at the edge of stability in excitatory-inhibitory neural populations
+Excitation-inhibition balance controls information encoding in neural populations
 Abstract
 Understanding how the complex connectivity structure of the brain shapes its information-processing capabilities is a long-standing question.
-Here, by focusing on a paradigmatic architecture, we study how the neural activity of excitatory and inhibitory populations encodes information on external signals.
-We show that information is maximized at the edge of stability, where excitation is balanced by inhibition.
-When the input switches among different stimuli, this maximum corresponds to the entropy of the external switching dynamics.
-By analyzing the case of a prolonged stimulus, we find that stronger inhibition is needed to maximize the instantaneous sensitivity, revealing an intrinsic trade-off between short-time responses and long-time accuracy.
-In agreement with recent experimental findings, our results open the avenue for a complete information-theoretic understanding of how and why inhibition strength should be tuned to optimize information-processing capabilities.
+By focusing on a paradigmatic architecture, we study how the neural activity of excitatory and inhibitory populations encodes information on external signals.
+We show that at long times information is maximized at the edge of stability, where inhibition balances excitation, both in linear and nonlinear regimes.
+In the presence of multiple external signals, this maximum corresponds to the entropy of the input dynamics.
+By analyzing the case of a prolonged stimulus, we find that stronger inhibition is instead needed to maximize the instantaneous sensitivity, revealing an intrinsic trade-off between short-time responses and long-time accuracy.
+In agreement with recent experimental findings, our results pave the way for a deeper information-theoretic understanding of how the balance between excitation and inhibitions controls optimal information-processing in neural populations.
 preprint: APS/123-QED
 From sensory perception to task-driven behaviors and decision-making processes, the brain constantly receives and integrates large amounts of environmental information.
-Both the encoding and processing of this information in the cortex involve a complex interplay among different neuronal populations koch1994large ; sakurai1996population ; quiroga2007decoding ; vyas2020computation .
-Understanding the basic mechanisms behind such information processing is thus a central topic in systems neuroscience.
-Several studies have investigated neural encoding at the level of individual neurons, showing that certain neurons selectively respond to specific features of incoming stimuli, such as spatial or temporal frequency, orientation, position, or depth kriegeskorte2021neural ; barlow1967neural ; campbell1968angular ; henry1974orientation , leading to a sparse encoding olshausen1996emergence .
-However, due to advancements in the ability to simultaneously record the activity of large numbers of neurons across various brain areas, recent decades have witnessed a shift of research focus towards the investigation of collective dynamics of neural populations gao2015simplicity ; buzsaki2004large ; brown2004multiple .
+Both the encoding and processing of information in the cortex involve a complex interplay among different neuronal populations koch1994large ; sakurai1996population ; quiroga2007decoding ; vyas2020computation whose understanding is a central topic in systems neuroscience.
+Several studies have investigated neural encoding at the level of individual neurons, showing that certain neurons selectively respond to specific features of incoming stimuli, such as spatial or temporal frequency, orientation, position, or depth kriegeskorte2021neural ; barlow1967neural ; campbell1968angular ; henry1974orientation ; olshausen1996emergence .
+Due to advancements in the ability to simultaneously record the activity of large numbers of neurons across brain areas, recent decades have witnessed a shift of research focus towards the investigation of collective dynamics of neural populations gao2015simplicity ; buzsaki2004large ; brown2004multiple .
 Remarkably, the trajectories of such populations are typically constrained in low-dimensional manifolds in the high-dimensional space of neural activity kriegeskorte2021neural ; buonomano2009state ; pandarinath2018inferring , suggesting that the entire population dynamically encodes stimulus variables in this reduced cunningham2014dimensionality ; mante2013context ; gallego2017neural ; remington2018flexible or coarse-grained meshulam2019coarse ; nicoletti2020scaling neural state space.
-Tools from information theory have been used to measure the amount of information that the response of a neural system conveys on a stimulus brunel1998mutual ; quian2009extracting ; bernardi2015frequency ; panzeri2022structures , for instance through mutual or Fisher information shannon1948mathematical ; blahut1987principles .
+Tools from information theory have been used to measure the amount of information that the response of a neural system conveys on a stimulus brunel1998mutual ; quian2009extracting ; bernardi2015frequency ; panzeri2022structures ; mariani2022disentangling , for instance through mutual or Fisher information shannon1948mathematical ; blahut1987principles .
 Yet, understanding how the emergent information properties depend on the underlying dynamics of the neural populations remains an open question.
-Recurrent neural networks trained on cognitive tasks have shown that low-dimensional subspaces naturally emerge to support flexible computations at the population level dubreuil2022role ; beiran2023parametric .
-However, the internal connectivities of these artificial networks do not adhere to biological constraints such as Dale’s principle, which states that neurons must be either excitatory or inhibitory.
-This neuro-anatomical feature has profound implications for functionality, as it underpins amplification mechanisms murphy2009balanced ; hennequin2014optimal ; apicella2021emergence ; schaub2015emergence and efficient transmission baggio2020efficient .
-It has been recently shown that strong recurrent coupling and inhibition stabilization are common features of the cortex sanzeni2020inhibition .
+Strong recurrent coupling and inhibition stabilization are common features of the cortex sanzeni2020inhibition .
 Crucially, such a finely balanced state not only prevents instability but may enhance the system’s computational properties as well.
-Networks operating in this near-critical state often perform better in information processing tasks and complex computations in general langton1990computation ; bertschinger2004real ; beggs2008criticality , while exhibiting optimal sensitivity to sensory stimuli kinouchi2006optimal ; shew2009neuronal .
+Networks operating in this fine-tuned state often perform better in information processing tasks and complex computations langton1990computation ; bertschinger2004real ; beggs2008criticality , while exhibiting optimal sensitivity to sensory stimuli kinouchi2006optimal ; shew2009neuronal .
 Such an interplay between strong excitatory coupling and compensatory inhibition is shaped by the connectivity structure between neural populations, which makes theoretical studies particularly challenging.
 In this work, we explicitly tackle the problem of quantifying the information encoded by neuronal subpopulations on an external stochastic stimulus.
-By leveraging neural dynamical models and focusing on the connectivity between excitatory and inhibitory populations, we are able to compute the information between the neuronal system and the stimulus in suitable regimes and provide analytical bounds on the mutual information.
-Specifically, we demonstrate that an excitatory-inhibitory balanced state is necessary not only to ensure stability but also for maximizing information encoding at the steady state.
-Further, by studying the response of the system to a single stochastic perturbation of varying intensity, we reveal an intrinsic trade-off between the optimal response at short and long times.
-In particular, the time-resolved form of the mutual information shows that global inhibition acts to regulate the total information encoded and the sensitivity of the system’s response.
-To retain physical interpretability, we focus on a coarse-grained model of two neuronal subpopulations, one excitatory and one inhibitory, described by their respective firing rates $\rho_{E}$ and $\rho_{I}$.
-The excitatory population receives a time-varying external input $h(t)$, representing the stimuli the neurons seek to encode in their firing rate dynamics.
-We are interested in the fluctuations around the steady states of the two populations, $\rho_{\mu}^{*}$, which we define as $x_{\mu}(t)=\rho_{\mu}(t)-\rho_{\mu}^{*}$ with $\mu\in\{E,I\}$.
-This model has been widely used in the literature as it captures essential properties of neuronal connectivity murphy2009balanced ; schaub2015emergence ; christodoulou2022regimes ; wu2023inhibition .
-Following these previous works, we obtain the Langevin equations:
-$\tau\frac{dx_{\mu}}{dt}=-r_{\mu}x_{\mu}+\sum_{\nu\in{E,I}}A_{\mu\nu}x_{\nu}+\sqrt{2D_{\mu}}\xi_{\mu}+h(t)\delta_{\mu,E}$ (1)
-where $\tau$ is the characteristic timescale of neural dynamics, $r_{\mu}$ is the decay term of the activity, $\xi_{\mu}$ are independent white noises, and $A_{\mu\nu}$ is an element of the synaptic connectivity matrix, that is parametrized as
+By focusing on the connectivity between excitatory and inhibitory populations, we compute the information between the neural activity and the external stimuli, providing analytical bounds on the mutual information in suitable limits.
+We demonstrate that an excitatory-inhibitory balanced state is necessary not only to ensure stability but also to maximize information encoding at the steady state, both in linear and nonlinear regimes.
+Further, by studying the response to a single stochastic perturbation of varying intensity, we reveal an intrinsic trade-off between the optimal response at short and long times.
+In particular, global inhibition acts to regulate the total information encoded and the sensitivity of the system’s response.
+To retain physical interpretability, we consider the activity of two neuronal subpopulations, one excitatory $x_{E}$, and one inhibitory $x_{I}$.
+The excitatory population receives a time-varying external input $h(t)$, representing the stimuli the neurons seek to encode in their dynamics, described by the Langevin equations
+$\tau\frac{dx_{\mu}}{dt}=$ $-r_{\mu}x_{\mu}+\sum_{\nu\in{E,I}}A_{\mu\nu}f(x_{\nu})+$ (1)
+$+h(t)\delta_{\mu,E}+\sqrt{2D_{\mu}\tau}\xi_{\mu}$
+where $\tau$ is the characteristic neural timescale, $r_{\mu}$ is the decay of the activity, $\xi_{\mu}$ are independent white noises, and $f$ is an activation function.
+$A_{\mu\nu}$ is an element of the synaptic connectivity matrix
 $\hat{A}=\begin{pmatrix}w&-kw\\ w&-kw\end{pmatrix}\,,$ (2)
@@
 Thus, $w$ measures the overall excitation strength, while $k$ quantifies the relative intensity of the inhibition.
-One can immediately show that, if $r_{\mu}=r$, the system is stable when $k>k_{c}=1-r/w$, while for $k<k_{c}$ the excitation is too strong for the inhibition to stabilize it.
-To model changes in the external environment, we posit the input as a jump process between a ground state, $h_{0}$, and a set of $M$ discrete environmental states, $h_{i}$.
-These states might represent sensory stimuli of different intensities, behavioral states, or motor commands.
-For simplicity, we take these states to be equispaced, such that $h_{i}=h_{0}+i\Delta h$, with $h_{0}=0$ representing the absence of external signals.
-The environment can transition from this ground state to any other state with uniform transition rates, which we denote by $w_{0\to i}=w_{\uparrow}$.
-Likewise, the transition rate to the ground state from any other state is $w_{i\to 0}=w_{\downarrow}$.
-All other transition rates are set to zero, i.e., there is no direct switch from one stimulus to another.
-In this setting, all environmental states are equally likely and the neural populations must respond to the stochastic jumps among them.
-The characteristic timescale of this environmental process is $\tau_{\mathrm{input}}=(w_{\uparrow}+w_{\downarrow})^{-1}$.
-We sketch the model in Fig.
-1a.
-Figure 1: (a) Sketch of the model, describing a population of excitatory ($E$, green) and inhibitory neurons ($I$, blue) in a linearized regime evolving on a timescale $\tau$.
-An input $h$ stimulates activity in the excitatory population with a timescale $\tau_{\mathrm{input}}$.
-(b-c) If $\tau_{\mathrm{input}}\ll\tau$, the neural populations are not able to resolve the different inputs.
+This model (Fig.
+1a) has been widely used in the literature as it captures essential properties of neuronal connectivity murphy2009balanced ; schaub2015emergence ; christodoulou2022regimes .
+In the Supplemental Material (SM) suppinfo_ref , we also study the effect of different input projections by considering the case in which both populations receive the input.
+To model changes in the external environment, the input follows a jump process between a ground state $h_{0}$ and a set of $M$ environmental states $h_{i}$ describing, e.g., sensory stimuli of different intensities, behavioral states, or motor commands.
+For simplicity, we take $h_{i}=h_{0}+i\Delta h$, with $h_{0}=0$ representing the absence of external signals and $\Delta h$ is a constant shift between the inputs.
+The input switches from $h_{0}$ to any $h_{i}$ with uniform transition rates $q_{0\to i}=q_{\uparrow}$, and similarly $q_{i\to 0}=q_{\downarrow}$.
+All other transition rates are set to zero, i.e., no direct switches from one stimulus to another are present.
+Hence, all environmental states are equally likely and the neural populations must respond to the stochastic jumps among them.
+The characteristic timescale of the input process is $\tau_{\mathrm{input}}=(q_{\uparrow}+q_{\downarrow})^{-1}$.
+Figure 1: (a) Sketch of the model, describing a population of excitatory ($E$, green) and inhibitory neurons ($I$, blue) evolving on a timescale $\tau$.
+An input $h$ stimulates activity in the excitatory population on a timescale $\tau_{\mathrm{input}}$.
+(b-c) In this figure, $f(x)=x$.
+If $\tau_{\mathrm{input}}\ll\tau$, neural populations are not able to resolve different inputs.
 In the opposite limit, the joint probability $p(E,I)$ displays instead peaks around the different input strengths.
-(d) The mutual information $I_{\mathbf{x},h}$ is zero in the fast-inputs regime, but sharply increases when $\tau_{\mathrm{input}}\gg\tau$, signaling that the neural populations are capturing information on the input in this regime.
-Simulations obtained with $M=2$, $\tilde{w}_{\uparrow}=1/3,\tilde{w}_{\downarrow}=2/3,D=1/2,r=1,\tau=1$.
+(d) $I_{\mathbf{x},h}$ is evaluated by sampling numerically Eq.
+(1).
+The mutual information is zero in the fast-inputs regime, but sharply increases when $\tau_{\mathrm{input}}\gg\tau$, signaling that neural populations are capturing information on the input.
+Parameters: $M=2$, $\tilde{q}_{\uparrow}=1/3$, $\tilde{q}_{\downarrow}=2/3$, $D=1/2$, $r=1$, $\tau=1$, $w=2$, $k=1.1$, $\Delta h=2.5$ in all figures, unless stated otherwise.
+Analytical bounds are in Eq.
+(5).
 We seek to understand how much information the neuronal network can capture on the external inputs at stationarity.
@@
 $=H_{\bm{x}}+H_{\mathrm{input}}-H_{\bm{x},\mathrm{input}}$ (3)
-which is nothing but the Kullback-Leibler divergence $D_{\mathrm{KL}}$ between the joint steady-state probability of the inputs and the neural activity, $p^{\mathrm{st}}_{i,\bm{x}}$, with $\bm{x}=(x_{E},x_{I})$, and the corresponding marginal distributions, $p^{\mathrm{st}}_{\bm{x}}$ and $\pi_{i}^{\mathrm{st}}$ cover1999elements .
-As such, $I_{\bm{x},h}^{\mathrm{st}}$ quantifies how much $\bm{x}$ and $h$ are dependent on one another in terms of how much information they share in their joint probability distribution.
-As in Eq.
-(3), the mutual information can be rewritten in terms of the differential entropy of the excitatory and inhibitory populations, $H_{\bm{x}}$, the Shannon entropy of the external inputs, $H_{\mathrm{input}}$, and the joint entropy of the neural activity and environment, $H_{\bm{x},\mathrm{input}}$.
-Without loss of generality, we now assume that the populations have the same diffusion coefficient, i.e., $D_{\mu}=D$, and the same decay rate, i.e., $r_{\mu}=r$.
-From Eq. (1), the joint probability $p_{i,\bm{x}}(t)$ is governed by the following Fokker-Planck equation:
+where $H_{\bm{x}}$ is the differential entropy of the excitatory and inhibitory populations, $H_{\mathrm{input}}$ the Shannon entropy of the external inputs, and $H_{\bm{x},\mathrm{input}}$ their joint entropy.
+Eq.
+(3) can be understood as the Kullback-Leibler divergence between the joint steady-state probability of the inputs and the neural activity, $p^{\mathrm{st}}_{i,\bm{x}}$, with $\bm{x}=(x_{E},x_{I})$, and the corresponding marginal distributions, $p^{\mathrm{st}}_{\bm{x}}$ and $\pi_{i}^{\mathrm{st}}$ cover1999elements .
+As such, $I_{\bm{x},h}^{\mathrm{st}}$ quantifies all statistical dependencies between $\bm{x}$ and $h$ in terms of how much information is encoded in the joint probability of the input and the neural activity.
+For simplicity, we take $D_{\mu}=D$ and $r_{\mu}=r$.
+From Eq. (1), the joint probability $p_{i,\bm{x}}(t)$ is governed by the Fokker-Planck equation:
 $\partial_{t}p_{i,\bm{x}}(t)=$ $\dfrac{1}{\tau}\sum_{\mu=E,I}\biggl{[}\partial_{\mu}[(\tilde{F}_{i\mu}(\bm{x})p_{i,\bm{x}}(t)]+\partial_{\mu}^{2}p_{i,\bm{x}}(t)\biggl{]}+$
-$+\dfrac{1}{\tau_{\mathrm{input}}}\sum_{j=0}^{M}\biggl{[}\tilde{w}_{j\to i}p_{j,\bm{x}}(t)-\tilde{w}_{i\to j}p_{i,\bm{x}}(t)\biggl{]}$ (4)
-where $\tilde{w}_{j\to i}=\tau_{\mathrm{input}}w_{j\to i}$ are the rescaled transition rates, $\tilde{F}_{i\mu}(\bm{x})=-rx_{\mu}+\sum_{\nu}\mathcal{A}_{\mu\nu}x_{\nu}+h_{i(t)}\delta_{\mu,E}$ with $i(t)$ the environmental state at time $t$, and we used the shorthand notation $\partial_{x_{\mu}}:=\partial_{\mu}$.
+$+\dfrac{1}{\tau_{\mathrm{input}}}\sum_{j=0}^{M}\biggl{[}\tilde{q}_{j\to i}p_{j,\bm{x}}(t)-\tilde{q}_{i\to j}p_{i,\bm{x}}(t)\biggl{]}$ (4)
+where $\tilde{q}_{j\to i}=\tau_{\mathrm{input}}q_{j\to i}$ are the rescaled transition rates, $\tilde{F}_{i\mu}(\bm{x})=-rx_{\mu}+\sum_{\nu}A_{\mu\nu}f(x_{\nu})+h_{i(t)}\delta_{\mu,E}$ with $i(t)$ the environmental state at time $t$, and we used the shorthand notation $\partial_{x_{\mu}}:=\partial_{\mu}$.
 Finding an explicit solution of Eq.
@@
 However, exact solutions can be accomplished in a timescale separation limit nicoletti2021mutual ; nicoletti2022mutual ; nicoletti2022information .
-We find that, in the limit of a fast-evolving input $\tau_{\mathrm{input}}\ll\tau$, the joint probability of the system factorizes as $p_{i,\bm{x}}^{\mathrm{st}}=p_{\bm{x}}^{\mathrm{st}}\pi_{i}^{\mathrm{st}}$ (see Supplemental Material suppinfo_ref ).
-In particular, $p_{\bm{x}}^{\mathrm{st}}$ corresponds to the solution of the Langevin equations in Eq.
-(1) with an effective input $\tilde{h}_{\mu}=\delta_{\mu,E}\sum_{i}h_{\mu,i}\pi_{i}^{\mathrm{st}}$ (Figure 1b).
-In this regime, we clearly have that the mutual information between the neural populations and the input vanishes, i.e., $I_{\bm{x},h}\to 0$ when $\tau_{\mathrm{input}}/\tau\to 0$.
-Indeed, the stationary solution of the system tells us that the neural activity is only influenced by the average input, as it cannot resolve its fast temporal evolution.
-On the other hand, in the limit of a slowly evolving external input $\tau_{\mathrm{input}}\gg\tau$, we find that the system is described by the stationary probability $p_{i,\bm{x}}(t)=p_{\bm{x}|i}^{\mathrm{st}}\pi_{i}(t)$, where $p^{\mathrm{st}}_{\bm{x}|i}$ is the probability of the excitatory and inhibitory populations conditioned on a given environmental state $h_{i}$.
-In particular, $p^{\mathrm{st}}_{\bm{x}|i}$ is the stationary solution of the system of Langevin equations in Eq.
-(1) with constant input $h_{i}$, which is a multivariate Gaussian distribution $\mathcal{N}(\bm{m}^{\mathrm{st}}_{i},\hat{\Sigma}^{\mathrm{st}})$.
-The mean $\bm{m}^{\mathrm{st}}_{i}$ is given by
+In the limit of a fast-evolving input $\tau_{\mathrm{input}}\ll\tau$, the joint probability of the system factorizes as $p_{i,\bm{x}}^{\mathrm{st}}=p_{\bm{x}}^{\mathrm{st}}\pi_{i}^{\mathrm{st}}$ (see SM suppinfo_ref ).
+$\pi_{i}^{\mathrm{st}}$ is the stationary distribution of the inputs, and $p_{\bm{x}}^{\mathrm{st}}$ is the solution of Eq.
+(1) with an effective input $\tilde{h}_{\mu}=\delta_{\mu,E}\sum_{i}h_{\mu,i}\pi_{i}^{\mathrm{st}}$ (Fig.
+1b).
+In this regime, the mutual information between the neural populations and the input vanishes, i.e., $I_{\bm{x},h}\to 0$ when $\tau_{\mathrm{input}}/\tau\to 0$.
+Indeed, the stationary solution of the system shows that the neural activity is only influenced by the average input, as it cannot resolve its fast temporal evolution.
+On the other hand, in the limit of a slowly evolving external input $\tau_{\mathrm{input}}\gg\tau$, the system is described by the stationary probability $p_{i,\bm{x}}(t)=p_{\bm{x}|i}^{\mathrm{st}}\pi_{i}(t)$, where $p^{\mathrm{st}}_{\bm{x}|i}$ is the probability of the excitatory and inhibitory populations at constant input $h_{i}$.
+We first focus on a linear activation function $f(x)=x$.
+In this case, the system is stable when $k>k_{c}=1-r/w$, while excitation is too strong for the inhibition to stabilize it for $k<k_{c}$.
+Then, $p^{\mathrm{st}}_{\bm{x}|i}$ is a multivariate Gaussian distribution $\mathcal{N}(\bm{m}^{\mathrm{st}}_{i},\hat{\Sigma}^{\mathrm{st}})$ with mean $\bm{m}^{\mathrm{st}}_{i}$
 $\begin{pmatrix}{m}^{\mathrm{st}}_{E,i}\\ {m}^{\mathrm{st}}_{I,i}\end{pmatrix}=\left(\hat{R}-\hat{A}\right)^{-1}\begin{pmatrix}h_{i}\\ 0\end{pmatrix}$
-where $R_{\mu\nu}=r_{\mu}\delta_{\mu\nu}$, whereas the covariance $\hat{\Sigma}^{\mathrm{st}}$ satisfies the Lyapunov equation:
-$\sum_{\nu\in\{E,I\}}\left[\left(A_{\alpha\nu}-R_{\alpha\nu}\right)\Sigma^{\mathrm{st}}_{\nu\beta}+\Sigma^{\mathrm{st}}_{\alpha\nu}\left(A_{\nu\beta}^{T}-R_{\nu\beta}\right)\right]=-\delta_{\alpha\beta}$
-as we show in the Supplemental Material suppinfo_ref .
+where $R_{\mu\nu}=r_{\mu}\delta_{\mu\nu}$, and a covariance $\hat{\Sigma}^{\mathrm{st}}$ that satisfies the Lyapunov equation:
+$\sum_{\nu}\left[\left(A_{\alpha\nu}-R_{\alpha\nu}\right)\Sigma^{\mathrm{st}}_{\nu\beta}+\Sigma^{\mathrm{st}}_{\alpha\nu}\left(A_{\nu\beta}^{T}-R_{\nu\beta}\right)\right]=-{\color[rgb]{0,0,0}2D_{\alpha}}\delta_{\alpha\beta}$
+as we show in the SM suppinfo_ref .
 It is worth noting here that, since the input acts as an additional drift, it only changes the average of the distribution with respect to the case of no input.
 Therefore, the stationary probability distribution of the neural populations is the Gaussian mixture $p_{\bm{x}}^{\mathrm{st}}=\sum_{i}\pi_{i}^{\mathrm{st}}p_{\bm{x}|i}^{\mathrm{st}}$.
-We show a typical trajectory of the system and the corresponding probability distribution of neural activity in Figure 1c.
-Even though the entropy of a Gaussian mixture cannot be written in a closed form, by employing the bounds proposed in kolchinsky2017estimating , we obtain an upper and a lower bound on the mutual information starting from the Chernoff-$\alpha$ divergence and the Kullback-Leibler divergence between the mixture components (see Supplemental Material suppinfo_ref ).
+We show a typical trajectory of the system and the corresponding probability distribution of neural activity in Fig.
+1c.
+Even though the entropy of a Gaussian mixture cannot be written in a closed form, by employing the bounds proposed in kolchinsky2017estimating , we obtain an upper and a lower bound on the mutual information starting from the Chernoff-$\alpha$ divergence and the Kullback-Leibler divergence between the mixture components (see SM suppinfo_ref ).
 We have that $I^{(b)}(\eta/4)\leq I_{\bm{x},h}\leq I^{(b)}(\eta)$, where
-$I^{(b)}(\eta)$ $=-\sum_{i=0}^{M}\pi_{i}\log[\sum_{j=0}^{M}\pi_{j}e^{-(j-i)^{2}\eta}\biggl{]}$ (5)
+$I^{(b)}(\eta)$ $=-\sum_{i=0}^{M}\pi_{i}^{\mathrm{st}}\log[\sum_{j=0}^{M}\pi_{j}^{\mathrm{st}}e^{-(j-i)^{2}\eta}\biggl{]}$ (5)
 with
-$\eta=\frac{\Delta h^{2}}{4Dr}\dfrac{(2r+(k-1)w)(2r^{2}+(3k-1)w+(k^{2}+1)w^{2})}{(r+(k-1)w)(2r^{2}+2(k-1)rw+(k^{2}+1)w^{2})}\;.$
+$\eta=\frac{\Delta h^{2}}{4Dr}\dfrac{[r+w(k-k_{c})][2r^{2}+(3k-1)w+(k^{2}+1)w^{2}]}{w^{2}(k-k_{c})(2r(k-k_{c})+(k^{2}+1)w)}\;.$
 Since $I^{(b)}(\eta/4)>0$, we have that $I_{\bm{x},h}$ is always non-zero.
@@
 In the intermediate regime between the fast- and slow-input limits, we cannot solve the Fokker-Planck equation explicitly.
-However, a direct simulation of the system shows that the mutual information smoothly interpolates between the two regimes, as we see in Figure 1d.
-Taken together, our results underscore the significance of timescales for neuronal circuits and their capability of processing information on external time-varying stimuli butts2007temporal ; das2019critical ; mariani2021b ; nicoletti2024information .
-Figure 2: (a) Mutual information between the neural activity $\bm{x}=(x_{E},x_{I})$ and the input $h$, as a function of the excitation strength $w$ and the relative intensity of inhibition $k$.
-If $k<k_{c}$ ($k_{c}$ is the black dotted line), the system is unstable.
-(b-c) Information between the neural activity and the inputs is maximized at the edge of stability when $k\to k_{c}$.
-In this limit, $I_{\bm{x},h}$ converges to the entropy of the input $H_{\mathrm{input}}$, which quantifies the information contained in the input.
-Figure 3: (a) Dynamics of the mutual information in the presence of a prolonged stochastic input.
-(b-c) The stationary mutual information $I_{\bm{x},h}^{\mathrm{st}}$ diverges at the edge of stability, whereas the sensitivity $\chi_{\bm{x},h}=\partial_{t}^{2}I_{\bm{x},h}(t)|_{t=t_{\mathrm{stim}}}$ peaks at intermediate values of $k$ for large $w$.
-(d-e) The peak of the sensitivity occurs for $k_{\mathrm{max}}(w)>k_{c}$, while $I_{\bm{x},h}^{\mathrm{st}}\to+\infty$ for $k\to k_{c}$.
-(f) Thus, a larger inhibition strength $k$ benefits the response at short times in terms of how quickly information increases when the input signal arrives.
+However, a direct simulation of the system shows that the mutual information smoothly interpolates between the two regimes, as we see in Fig.
+1d.
+Taken together, our results underscore the significance of timescales for neuronal circuits and their capability of processing information on external time-varying stimuli butts2007temporal ; das2019critical ; mariani2021b ; nicoletti2024information ; nicoletti2024gaussian .
+Figure 2: (a-b) Mutual information between the neural populations and the input in the linear case.
+If $k<k_{c}$ (black dotted line), the system is unstable.
+Information is maximized at the edge of stability, $k\to k_{c}$.
+In this limit, $I_{\bm{x},h}$ converges to $H_{\mathrm{input}}$, which quantifies the information contained in the input.
+Mutual information is obtained from numerically integrating the Gaussian mixture, while analytical bounds are in Eq.
+(5).
+(c-d) Comparison with the nonlinear case.
+For small $\Delta h$, in the linear stability regime, results are comparable, whereas they differ for larger $\Delta h$.
+$I_{\bm{x},h}$ peaks at $k\approx k_{c}$, and decreases when $k<k_{c}$.
+Mutual information is obtained numerically with a kNN estimator.
+In (c), $D=0.001$, while $D=0.05$ in (d) to improve the numerical estimations.
+Figure 3: (a) Dynamics of mutual information with a constant stochastic input ($\sigma_{h}=1$).
+(b-c) $I_{\bm{x},h}^{\mathrm{st}}$ (Eq.
+(8)) diverges at the edge of stability, whereas $\chi_{\bm{x},h}$ (Eq.
+(9)) peaks at intermediate values of $k$ for large $w$.
+(d-e) Sensitivity peak occurs for $k_{\mathrm{max}}(w)>k_{c}$, while $I_{\bm{x},h}^{\mathrm{st}}\to+\infty$ for $k\to k_{c}$.
+(f) Thus, a larger inhibition strength $k$ benefits the short-time response when the input arrives.
 On the contrary, at long times, information is maximized by reducing $k$ and approaching the edge of stability at $k=k_{c}$.
 Crucially, the synaptic strengths of the excitatory and inhibitory populations drastically affect their mutual information with the input.
-Indeed, as we show in Figure 2a, $I_{\bm{x},h}$ strongly depends on the interplay between excitation and inhibition.
+Indeed, as we show in Fig.
+2a, $I_{\bm{x},h}$ strongly depends on the interplay between excitation and inhibition.
 Furthermore, the bounds in Eq.
@@
 2b):
-$I_{\bm{x},h}\underset{k\to k_{c}}{\longrightarrow}H_{\mathrm{input}}\;.$ (6)
+$I_{\bm{x},h}\underset{k\to k_{c}}{\longrightarrow}H_{\mathrm{input}}=-\sum_{i=0}^{M}\pi_{i}^{\mathrm{st}}\log\pi_{i}^{\mathrm{st}}\;.$ (6)
 Eq.
 (6) tells us that, at the edge of stability, the neural populations are able to fully capture the information contained in the external input, which is exactly its entropy $H_{\mathrm{input}}$.
-Intriguingly, we also find that this is the maximum value the mutual information can attain, as shown in Figure 2b-c.
+Intriguingly, we also find that this is the maximum value the mutual information can attain, as shown in Fig.
+2b.
 Thus, modulation of the inhibition strength plays a prime role in determining how efficiently the system can encode the external inputs, and the corresponding mutual information sharply increases as the edge of stability is approached.
+Notably, $\eta$ diverges also when $\Delta h^{2}/D\to\infty$, leading to maximal mutual information $H_{\mathrm{input}}$.
+This shows how the information encoded in neural activity crucially depends on the relative strength of input and noise (see SM suppinfo_ref ).
+Then, we consider a more realistic scenario of a nonlinear activation function $f(x)=\tanh(x)$.
+Since no analytic expression exists for $p_{\bm{x}|i}^{\mathrm{st}}$, we estimate $I_{\bm{x},h}$ numerically from the Langevin trajectories using a kNN estimator kraskov2004estimating .
+In Fig.
+2c, we show that for small $\Delta h$ the results are similar to the linear case in the stable region, whereas for larger $\Delta h$ we find quantitative differences between the two due to saturation effects (see Fig.
+2d and SM suppinfo_ref ).
+Crucially, in the presence of a nonlinear activation function the mutual information peaks at $k\approx k_{c}$ and decreases beyond the region of linear stability.
+These observations hint at the robustness of our results outside the linear scenario.
 So far, we have considered the steady-state response of the system to a time-varying input.
-However, understanding how neural populations dynamically acquire information when they receive a single stimulus, and how quickly they can do so, is crucial as well.
-To this end, we now introduce a scenario where the system in its stationary state is perturbed by an external stochastic input.
+However, understanding how and how quickly neural populations dynamically acquire information when a single persistent input is presented is crucial as well.
+We now consider a system in its stationary state that is perturbed by an external stochastic input.
 In this case, the mutual information in time reads
@@
 We assume that $p_{h}^{\mathrm{st}}$ is fully characterized by the mean input strength, $\mu_{h}$, and its variance, $\sigma_{h}$, so that $p_{h}=\mathcal{N}(\mu_{h},\sigma_{h})$.
-Then, we have that
-$I_{\bm{x},h}(t)$ $=\dfrac{1}{2}\log\dfrac{\det\left[\Sigma^{\mathrm{st}}+K(t)\begin{pmatrix}\sigma_{h}^{2}&0\\ 0&0\end{pmatrix}K(t)^{T}\right]}{\det\left(\Sigma^{\mathrm{st}}\right)}$ (8)
-where $\Sigma^{\mathrm{st}}$ is the covariance matrix of the unperturbed system at steady state, and $K(t)$ is the time-dependent gain matrix that we derive in the Supplemental Material suppinfo_ref .
-In Figure 3a, we plot the time evolution of the mutual information.
-At long times, we find once more that the mutual information is maximized at the edge of stability (Figure 3b), with $I_{\bm{x},h}$ diverging as $k\to k_{c}$.
+Then, we have:
+$I_{\bm{x},h}(t)=\dfrac{1}{2}\log\dfrac{\det\left[\hat{\Sigma}^{\mathrm{st}}+\hat{K}(t)\begin{pmatrix}\sigma_{h}^{2}&0\\ 0&0\end{pmatrix}\hat{K}(t)^{T}\right]}{\det\left(\hat{\Sigma}^{\mathrm{st}}\right)}$ (8)
+where $\hat{K}(t)$ is the time-dependent gain matrix that we derive in the SM suppinfo_ref .
+In Fig.
+3a, we plot the time evolution of the mutual information.
+At long times, we find once more that the mutual information is maximized at the edge of stability (Fig.
+3b), with $I_{\bm{x},h}$ diverging as $k\to k_{c}$.
 We note that, since the differential entropy for the continuous input distribution is not necessarily positive, the bounds in Eq.
@@
 In the inhibition-stabilized regime, where $w>1$, the response of the neural populations exhibits a faster increase for stronger excitatory couplings away from the edge of instability.
-To assess the system’s responsiveness, we introduce a metric of sensitivity defined as $\chi_{\bm{x},h}=\partial_{t}^{2}I_{\bm{x},h}(t)|_{t=t_{\mathrm{stim}}}$, which quantifies the instantaneous increase in mutual information when the stimulus arrives at time $t=t_{\mathrm{stim}}$ derivative_note .
-In Figure 3c, we show that $\chi_{\bm{x},h}$ peaks at an optimal inhibition strength $k_{\rm max}>1$, whose complete expression is given in suppinfo_ref .
-Crucially, this optimal value depends on the excitation strength as well, i.e., $k_{\rm max}\equiv k_{\rm max}(w)$ (see Fig.
+To assess the system’s responsiveness, we introduce a metric of sensitivity:
+$\chi_{\bm{x},h}$ $=\frac{\partial^{2}I_{\bm{x},h}(t)}{\partial t^{2}}\biggr{|}_{t=t_{\mathrm{stim}}}$ (9)
+$=\frac{\sigma_{h}^{2}}{2D}\frac{[r+w(k-k_{c})][2wk_{c}^{2}+r(k+1)]}{2r(k-k_{c})+w(1+k^{2})}$
+which measures how quickly the information on the stimulus increases immediately after the stimulation time $t=t_{\mathrm{stim}}$ derivative_note .
+In Fig.
+3c, we show that $\chi_{\bm{x},h}$ peaks at an optimal inhibition strength $k_{\rm max}(w)>1$, whose complete expression is given in the SM suppinfo_ref .
+Crucially, this optimal value depends on the excitation strength (see Fig.
 3d).
-This highlights that the inhibition regime for the optimal response at short times is drastically different from that at long times (see Figs.
-3d-e).
-In particular, since $k_{c}(w)<k_{\mathrm{max}}(w)$ for all $w$, our results unravel a fundamental trade-off between achieving maximum accuracy and the speed at which the neural populations encode information about the external stimulus, akin to a speed-accuracy trade-off lan2012energy .
-As we show in Figure 3f, long-time accuracy is generally achieved at lower inhibition, whereas sensitivity maximization requires a larger value of $k$.
+This reveals that the inhibition regime for the optimal response at short times is drastically different from that at long times, as we show in Figs.
+3d-e.
+In particular, since $k_{c}(w)<k_{\mathrm{max}}(w)$ for all $w$, our results unravel a fundamental trade-off between achieving maximum accuracy and the speed at which the neural populations encode information about the external stimulus, akin to a speed-accuracy trade-off emerging in different biological contexts lan2012energy .
+Remarkably, similar trade-offs have been recently found by studying the response of recurrent neural networks with random connectivities at different observation times azizpour2023available .
+As we show in Fig.
+3f, long-time accuracy is generally achieved at lower inhibition, whereas sensitivity maximization requires a larger value of $k$.
+In the SM suppinfo_ref , we show that, when both populations receive the prolonged stimulus, sensitivity is maximized at $k\to\infty$, suggesting that inhibition-dominated networks exhibit an enhanced tracking ability of the input renart2010asynchronous .
 Overall, our analysis marks a significant step towards the understanding of how the structure of connectivity shapes information encoding in neuronal population dynamics.
-We showed analytically in an exactly solvable regime that the mutual information between an input and the receiving neuronal populations is maximized at the edge of stability, highlighting the importance of the balance between excitation and inhibition.
-As non-normal synaptic interactions are crucial for achieving this optimal state, our findings confirm that the underlying structural connectivity is essential for the emergence of critical behaviors that support optimal information encoding and complex dynamical evolution both in whole-brain connectomes barzon2022criticality and in artificial recurrent neural networks kerg2019non .
+We have shown, both analytically in an exactly solvable regime and numerically for a nonlinear scenario, that the mutual information between a switching input and the receiving neuronal populations is controlled by the balance between the excitatory and inhibitory couplings, and peaks at the edge of linear stability.
+Moreover, we found that an increased inhibition strength is instrumental for establishing a robust response at short times, highlighting the importance of precisely tuning excitation and inhibition to achieve optimal encoding at different timescales.
+As non-normal synaptic interactions are crucial for realizing this optimal state, our findings at a coarse-grained level underscore the essential role of the underlying connectivity.
+In particular, structural connectivity has been shown to be essential in supporting complex dynamical evolution both in whole-brain connectomes barzon2022criticality and in artificial recurrent neural networks baggio2020efficient .
+Our study opens the avenue for an information-theoretic quantification of these emerging features from first principles, at the level of neural populations.
+In future works, the extension of similar ideas to more microscopic models might shed light on the intimate link between complex, even chaotic langton1990computation ; clark2024theory , dynamics, and information processing performances.
 Notably, alterations in excitatory-inhibitory balance have been experimentally related to the loss of information-processing efficiency observed in pathological conditions dehghani2016dynamic ; sohal2019excitation .
-Further, our results point towards the hypothesis that the bran might be poised at criticality to optimize its information-processing capabilities beggs2008criticality ; o2022critical ; mariani2021b ; mariani2022disentangling , and that computation at edge-of-chaos may be optimal langton1990computation .
 Our predictions are also consistent with recent experimental studies in which theoretical tools from response theory have been applied to extensive whole-brain neuronal recordings.
 The emergent dynamics of several brain regions has been shown to lie at the edge of stability dahmen2019second ; morales2023quasiuniversal , with a distance from instability that only slightly varied along the cortex.
-Such heterogeneity might be explained as an increase in the inhibition level wang2020macroscopic .
-Our findings suggest that this observed feature may be related to the tuning of sensitivity to different timescales murray2014hierarchy ; manea2022intrinsic .
-Importantly, the external stimulus considered here may be immediately generalized to a high-dimensional signal representing, for example, multiple stimuli with different characteristics (e.g., frequency, intensity) targeting spatially separated populations.
-Although we focused on a paradigmatic - yet widely used - model, our approach applies in principle to any neuronal connectivities and can be used to investigate more detailed synaptic structures and their effects on information-processing performances.
-Although it will be crucial to understand how strongly non-linear regimes affect the behavior of the mutual information and the information-theoretic features of the system in general, our work paves the way to the unraveling of the fundamental mechanisms supporting information encoding and sensitivity in neuronal networks.
+Such heterogeneity might be explained as an increase in the inhibition level wang2020macroscopic , and our findings suggest that this observed feature may be related to the tuning of sensitivity to different timescales murray2014hierarchy ; manea2022intrinsic .
+Importantly, the external input considered here may be immediately generalized to a high-dimensional signal representing, for example, multiple stimuli with different characteristics (e.g., frequency, intensity) targeting spatially separated populations potentially evolving on different timescales.
+Although we focused on a paradigmatic - yet widely used - model, our approach can be extended to investigate more detailed and microscopic synaptic structures and include the role played by plasticity to drive accurate encoding.
+Overall, our work paves the way to the unraveling of the fundamental mechanisms supporting information encoding and sensitivity in neuronal networks.
 Acknowledgements.
-G.N. acknowledges funding provided by the Swiss National Science Foundation through its Grant CRSII5_186422.
+Acknowledgments.—G.N. acknowledges funding provided by the Swiss National Science Foundation through its Grant CRSII5_186422.